### v2

# Explicit vs. Implicit: Investigating Social Bias in Large Language Models through Self-Reflection

Abstract
Large Language Models (LLMs) have been shown to exhibit various biases and stereotypes in their generated content. While extensive research has investigated bias in LLMs, prior work has predominantly focused on explicit bias, leaving the more nuanced implicit biases largely unexplored. This paper presents a systematic framework grounded in social psychology theories to investigate and compare explicit and implicit biases in LLMs. We propose a novel "self-reflection" based evaluation framework that operates in two phases: first measuring implicit bias through simulated psychological assessment methods, then evaluating explicit bias by prompting LLMs to analyze their own generated content. Through extensive experiments on state-of-the-art LLMs across multiple social dimensions, we demonstrate that LLMs exhibit a substantial inconsistency between explicit and implicit biases, where explicit biases manifest as mild stereotypes while implicit biases show strong stereotypes. Furthermore, we investigate the underlying factors contributing to this explicit-implicit bias inconsistency. Our experiments examine the effects of training data scale, model parameters, and alignment techniques. Results indicate that while explicit bias diminishes with increased training data and model size, implicit bias exhibits a contrasting upward trend. Notably, contemporary alignment methods (e.g., RLHF, DPO) effectively suppress explicit bias but show limited efficacy in mitigating implicit bias. These findings suggest that while scaling up models and alignment training can address explicit bias, the challenge of implicit bias requires novel approaches beyond current methodologies.
Explicit vs. Implicit: Investigating Social Bias in Large Language Models through Self-Reflection
Yachao Zhao, Bo Wangthanks: *Corresponding author., Yan Wang, College of Intelligence and Computing, Tianjin University, Tianjin, China {zhaoyachao, bo_wang}@tju.edu.cn
1 Introduction
Figure 1: Our proposed assessment methodology based on self-reflection. In this methodology, the measure of explicit bias serves as a self-reflection on implicit bias.
Research in social bias within Large Language Models (LLMs) has made substantial progress, with researchers developing various methods to identify and quantify model biases. However, current bias research faces significant limitations - these studies either fail to differentiate between explicit and implicit biases in LLMs, or predominantly focus on explicit bias while overlooking implicit bias. These limitations result in an incomplete understanding of model bias, highlighting the need for a more comprehensive research on the relation between explicit and implicit bias.
In social psychology, social bias is categorized into two forms, explicit and implicit biasGreenwald and Banaji (1995). Explicit bias, typically manifesting as conscious and controllable attitude expressions, is primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011). In contrast, implicit bias, which reflects unconscious and automated cognitive processes, is commonly measured by the Implicit Association Test (IAT) Greenwald et al. (1998). Extensive psychological research examining the relation between explicit and implicit biases in human individuals has revealed frequent inconsistencies between these two forms of bias, particularly regarding sensitive social targets such as gender and race. Furthermore, psychological research has extensively investigated the causes of this inconsistency, primarily attributing it to early learning processesBaron and Banaji (2006) and social expectationsCrandall et al. (2002). However, these investigations remain predominantly focused on human individuals, with limited research examining the explicit and implicit biases exhibited by large language models.
Based on these theoretical foundations, this study proposes an evaluation methodology based on prompt templates and self-reflection. This approach maps the measurement of explicit and implicit biases in LLMs to psychological assessment methods: SRA and IAT, respectively, thereby extending evaluation methods from human subjects to LLMs. The core of this methodology lies in measuring explicit bias through model self-reflection (Self-reflection) on its manifested implicit biases. It enables the comparative analysis of explicit versus implicit biases regarding identical targets.
To comprehensively evaluate and compare explicit and implicit biases in LLMs, this study conducted experiments on a representative selection of models, including both proprietary and open-source mainstream LLMs such as ChatGPTOuyang et al. (2022), Claude-3.5-SonnetBai et al. (2022b), and LLaMA-3.1-405B-InstructDubey et al. (2024). The experiments encompassed multiple social dimensions including gender, race, occupation, age, and disability. Results reveal a phenomenon that all LLMs exhibit bias inconsistency patterns similar to humans across different dimensions: demonstrating minimal stereotyping at the explicit level while manifesting significant stereotypical associations at the implicit level.
Given these findings, this study further investigates the underlying causes of bias inconsistency in LLMs. While social psychology has extensively explored similar inconsistency in human individuals, the mechanisms behind such patterns in LLMs remain unclear. This study systematically examines three key factors that potentially contribute to bias inconsistency in LLMs: training data scale, model size, and alignment training. Our investigation focuses on the LLaMA model familyTouvron et al. (2023); Dubey et al. (2024), which provides an ideal experimental setting due to its open-source nature, transparent training data documentation, and availability of multiple model sizes with both base and instruction-tuned versions. Our experimental results reveal complex relation between these factors and bias manifestation in LLMs. As training data scale and model parameters increase, explicit bias shows a consistent decrease, while implicit bias demonstrates the opposite trend, showing steady increase. Alignment training exhibits a distinct impact - while it significantly reduces explicit bias, implicit bias remains relatively stable within a fixed range regardless of training steps. These findings suggest that while recent advances in LLM development have successfully reduced explicit biases through scaling and alignment techniques, addressing implicit biases may require fundamentally different approaches.
In summary, the main contributions of this paper are as follows:
1. We propose a novel "self-reflection" evaluation methodology aligned with psychological theory. Through carefully designed prompt templates, we employ SRA to measure explicit bias and IAT to measure implicit bias. Moreover, the assessment of explicit bias involves the model’s self-reflection on its implicit biases.
2. Our experimental results reveal inconsistency between explicit and implicit biases in LLMs. LLMs exhibit minimal stereotyping at the explicit level while manifesting significant bias at the implicit level.
3. We present the first systematic investigation of key factors influencing explicit and implicit biases in LLMs. Through in-depth experimental analysis across three dimensions - training data scale, model size, and alignment training - we uncover the differential impact mechanisms these factors have on both types of bias.
2 Related Work
2.1 Bias Research in Social Psychology
In social psychology, research on attitudinal bias represents a crucial area of investigation. Bias is commonly defined as preferential views toward specific targets Smith and Noble (2014). Researchers distinguish between two forms of bias: explicit and implicit Greenwald and Banaji (1995). Explicit bias refers to consciously recognized and deliberately expressed biased attitudes, which are overt and conscious in nature, primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011). The Likert Scale Likert (1932) is one of the most widely used measurement instruments in SRA, evaluating participants’ attitudes through 5-point or 7-point rating scales. Implicit bias refers to attitudes that individuals find difficult to identify and are typically unaware of, manifesting through underlying associations or automatic responses. These biases are commonly measured using the Implicit Association Test (IAT) Greenwald et al. (1998). In a typical IAT experiment, researchers evaluate implicit attitudes by measuring participants’ response times to different concept pairings. Social psychologists have conducted extensive comparative studies on explicit and implicit biases in individuals Dovidio et al. (2001); Nosek (2007); Son Hing et al. (2008). These studies have revealed a significant finding: when addressing sensitive social topics such as gender and race, individuals often demonstrate notable discrepancies between their explicit and implicit biases. For instance, research has shown that while participants explicitly express support for gender equality, their implicit attitudes reveal negative perceptions toward female scientists Moss-Racusin et al. (2012). This divergence is primarily attributed to individuals’ socialization processes and social expectations. Implicit biases typically form during early childhood, while the development of cognitive abilities enables individuals to progressively suppress the explicit expression of these implicit biases, resulting in this inconsistencyBaron and Banaji (2006). Additionally, this inconsistency is further amplified when mainstream social values conflict with implicit individual attitudesCrandall et al. (2002).
2.2 Bias Analysis in Large Language Models
Prior work has extensively investigated bias in Large Language Models. Some research has explored explicit bias in LLMs, typically measuring bias by directly including specific target objects in prompts or templatesAbid et al. (2021); Kirk et al. (2021); Kotek et al. (2023). Abid et al. Abid et al. (2021) found that GPT-3 Brown et al. (2020) tends to generate text with violence when given Muslim-related prompts. Other studies examine models’ implicit biasesCaliskan et al. (2017); Zhao et al. (2024). Marked PersonasCheng et al. (2023), through analyzing distributional differences in model-generated words across social groups, revealed that apparently positive model outputs contained implicit negative perceptions of Black women. Bai et al.Bai et al. (2024) developed a methodology based on the Implicit Association Test (IAT) to quantify implicit biases embedded in LLMs.
Beyond these bias detection methods, studies have also focused on investigating the factors influencing social bias in LLMs. Inspired by the Scaling Laws Kaplan et al. (2020), these studies primarily examine the relation between bias and various factors such as model size and training data composition. Furthermore, with the widespread use of Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) in LLM training, investigations have emerged to understand how alignment methods affect the bias in LLMs. Ganguli et al. Ganguli et al. (2023) focused on two key variables: model scale and RLHF training steps, examining their impact on models’ Moral Self-Correction capabilities. Their large-scale experiments revealed that within certain ranges, both model scale and RLHF training steps showed significant positive correlations with models’ ability to self-debias. Through a series of experiments, Ali et al. Ali et al. (2024) found that increasing model size may lead to higher levels of bias.
3 Methodology
Large language models(LLMs) possess the capability to reflect on and evaluate their own outputs, providing a new perspective for studying their biasesMadaan et al. (2023); Weng et al. (2023); Shinn et al. (2023). The evaluation method proposed in this study fully utilizes the model’s reflective capabilities: first measuring the model’s implicit biases, then measuring its explicit biases through the model’s reflection on these implicit biases. This method effectively compares both implicit and explicit biases toward the same social targets within the model, revealing the relation between these two type of biases.
Specifically, the evaluation method introduced in this section draws inspiration from psychological measures including the Implicit Association Test (IAT) and Self-Report Assessment (SRA), which are used to measure implicit and explicit biases respectively. Through carefully designed prompt engineering, these methods are adapted for evaluating LLMs.
3.1 Implicit Bias Assessment Based on Psychological Theory
Research in social psychology suggests that the key to measuring implicit bias lies in conducting assessments without subjects’ awareness Greenwald and Banaji (1995). The Implicit Association Test (IAT) is an important measurement tool designed based on this principle: rather than directly inquiring about subjects’ attitudes toward social groups such as gender or race, it measures implicit bias by examining differences in response times when subjects associate target-related stimuli (such as names of different genders/races) with corresponding attributes. For example, if a subject’s response time is significantly shorter for "male-career, female-family" combinations compared to "female-career, male-family" combinations, this may indicate an implicit bias associating men with careers and women with family roles.
To measure implicit bias in LLMs, this study innovatively adapts the IAT to model evaluation through prompt engineering. Specifically, we propose a prompt template containing attribute words and masked targets:
<mask> is attrX as <mask> is attrY,   (1) where $mask$ represents the masked target, while $attr_{X}$ and $attr_{Y}$ represent a predefined pair of attributes. This template requires the model to make a choice among given candidate words, thereby revealing implicit bias. To illustrate the methodology, we examine implicit associations between racial categories (White-Black) and valence attributes (good-bad). The experimental procedure consists of the following steps:
  1. 1.
    Determine target group pairs $T_{A}$ and $T_{B}$, corresponding to White and Black racial categories, and their respective stimulus sets $S_{a}$ and $S_{b}$ (e.g., sets of White and Black names).
  2. 2.
    Define attribute word pairs $A_{X}$ and $A_{Y}$, representing ’good’ and ’bad’ respectively.
  3. 3.
    Randomly select N stimuli from $S_{a}$ and $S_{b}$ to form subsets $s_{a}$ and $s_{b}$, and randomly choose one attribute word each ($a_{X}$ and $a_{Y}$) from $A_{X}$ and $A_{Y}$.
  4. 4.
    Insert $a_{X}$ and $a_{Y}$ into the attribute positions of the template, and combine and shuffle $s_{a}$ and $s_{b}$ to create a candidate word set $s_{ab}$.
  5. 5.
    Design a prompt including the above elements with corresponding instructions, requiring the model to select two words from $s_{ab}$ to fill the $mask$ positions in the template.
  6. 6.
    To enhance measurement reliability, following recommendations on prompt template diversification from existing research Seshadri et al. (2022); Goldfarb-Tarrant et al. (2023), we implemented several optimization methods: First, we created 5 semantically similar but structurally distinct test templates through synonym substitution and word modification. To further mitigate order effects, we swapped the attribute pairs in these 5 templates, resulting in 10 templates.
Beyond racial bias, this study conducts comparative experiments on both explicit and implicit biases across multiple sensitive social dimensions including gender, occupation, age, and disability, ensuring comprehensive experimental coverage.
3.2 Explicit Bias Assessment Through Self-Reflection
To measure explicit bias in LLM, this study draws on Self-report Assessment (SRA) from social psychology, a classic technique for measuring explicit bias. Unlike the indirect measurement of implicit bias, explicit bias measurement requires subjects to directly express their attitudes and views toward specific social groups while being fully aware of the measurement purpose. The Likert Scale Likert (1932) is one of the most commonly used measurement tools in self-report assessment, evaluating subjects’ attitudes using 5-point or 7-point rating scales.
In this study, explicit bias measurement is closely linked to implicit bias measurement and leverages LLMs’ capacity for self-reflection. Specifically, this paper measures explicit bias by guiding language models to systematically reflect on and evaluate their potential attitudes demonstrated during implicit bias measurement. Because this process requires LLMs to analyze and evaluate their implicit biases shown in Section 3.1, we term this approach the "self-reflection" assessment method.
In the implementation process, the measurement template needs to explicitly specify the target social groups, thus requiring the replacement of $mask$ in template (1) from the implicit test with specific social group concept words, such as "men" and "women" for gender categories, or "Black" and "White" for racial categories. Specifically, this paper randomly selects words from the predefined social group word sets $T_{A}$ and $T_{B}$ to replace $mask$, while keeping the predefined attribute words $attribute_{X}$ and $attribute_{Y}$ unchanged, resulting in the following template sentence:
TargetA is attrX as TargetB is attrY,    (2)
Subsequently, the LLM is prompted to evaluate the template sentence based on a five-point Likert scale, with options including: strongly disagree, disagree, neutral, agree, strongly agree. The model should assess its level of agreement based on the specific content of the sentence, select the most appropriate option from these five choices, and provide reasoning for its selection.
To enhance measurement reliability and mitigate potential order effects, this study randomizes the five Likert scale options during each test rather than using a fixed order. This operation helps reduce potential interference from option ordering on model choices, thereby maximizing the stability and consistency of measurement results.
4 Experiments
4.1 Experimental Setup
This study systematically examined six types of typical social stereotypes to comprehensively assess bias in LLMs. Four categories are derived from classical IAT Greenwald et al. (1998) and WEAT Caliskan et al. (2017) studies, utilizing their validated target words, attribute words, and corresponding stimuli. These four categories include:
1. Gender stereotypes in career-family domains: The strength of association between male/female and career/family.
2. Racial stereotypes: The degree of association between White/Black people and positive/negative emotions.
3. Age stereotypes: The strength of association between young/old and positive/negative emotions.
4. Gender stereotypes in academic domains: The degree of association between male/female and science/arts.
Beyond these four categories, this study added two important research subjects:
5. Disability stereotypes: The association between able-bodied/disabled and positive/negative emotions. Data for this part of research is based on the official Implicit Association Test website [^1] https://implicit.harvard.edu/implicit/takeatest.html
6. Occupational gender stereotypes: The degree of association between male/female and male-dominated/female-dominated occupations . Given the prevalence and research value of occupational gender bias studies in natural language processing Smith et al. (2022); Watson et al. (2023), this study collected 10 pairs of semantically similar occupations with significant gender distribution differences from the U.S. Bureau of Labor Statistics website [^2] https://www.bls.gov/cps/cpsaat11.htm. Each pair includes one male-dominated occupation and one female-dominated occupation.
For both implicit and explicit bias tests, we conducted 20 independent experiments on each of the 10 test templates per research subject, totaling 200 experiments for each type of bias. Given that this study involved 9 large language models, 6 types of stereotypes, and 2 types of bias (implicit and explicit), the total number of experiments reached 21,600 (9 × 6 × 200 × 2), a scale sufficient to ensure statistical reliability of the results. Following the research findings of Wang et al. Wang et al. (2023), all model evaluation parameters are set with Temperature=0 to reduce output randomness and obtain more reliable assessment results.
4.2 Experimental Metrics
Stereotypes refer to fixed, simplified, and often biased views or beliefs about a particular group or individualDovidio (2010). These views are typically based on characteristics such as race, gender, age, and occupation while often disregarding individual differences Gilbert and Hixon (1991). Our study focuses on examining the stereotypes reflected in both explicit and implicit biases in LLMs, and defines corresponding quantitative metrics.
In measuring implicit bias in this study, a stereotype is considered present when the LLM establishes semantic associations between each element of a pair of attributes and its corresponding traditional stereotypical target; otherwise, it is considered absent. This strict definition criterion helps identify the LLM’s implicit biases. For example, when the LLM associates White stimuli (such as White names) with positive attributes and Black stimuli (such as Black names) with negative attributes, it is considered to exhibit stereotyping.
In measuring explicit bias, as this study employs a five-point Likert scaleLikert (1932) for evaluation, stereotype expression statements are constructed based on the implicit bias template sentences, associating each element of a pair of attributes with its traditional stereotypical target, and requiring the LLM to make judgments. When the model selects "agree" or "strongly agree" and provides corresponding reasons, it is considered to exhibit stereotyping. When the LLM selects "neutral," "disagree," or "strongly disagree," it is considered to show no stereotyping.
To quantify the experimental results, this paper defines the "Stereotypical Score" (SC):
$Stereotypical\ Score=\frac{n_{stereotype}}{N}$ (1)
where $n_{stereotype}$ represents the number of stereotype expressions occurring in either explicit or implicit tests, and $N$ is the total number of experiments. This metric ranges from $[0,1]$, with scores closer to $1$ indicating more severe stereotyping and scores closer to 00 indicating less stereotyping.
4.3 Baselines
This study first selects high-ranking large language models from the Chatbot Arena platform Chiang et al. (2024) as research subjects. Chatbot Arena is an open language model evaluation platform that employs human preference-based assessment through a paired comparison mechanism: when users pose questions, the platform randomly selects two anonymous language models to respond, and users vote based on the quality of the models’ responses, with model identities revealed only after voting is complete. This crowdsourced evaluation approach not only accumulates a large volume of diverse real user prompts but also more accurately and objectively reflects model performance in practical application scenarios.
Specifically, among mainstream commercial models, this paper selected the following representative models: GPT-4-turboHurst et al. (2024), GPT-4oHurst et al. (2024), Claude-3.5-SonnetBai et al. (2022b), Gemini-2.0-FlashTeam et al. (2023), and others. To enhance the comprehensiveness of this experiment, this paper also included LLMs with significant influence in the open-source community: LLaMA-2-70B-chatTouvron et al. (2023) and LLaMA-3.1-405B-InstructDubey et al. (2024). Table 1 presents detailed information about these selected LLMs.
Model Open Source Rank
Gemini-2.0-Flash $\times$ 1
Claude-3.5-Sonnet $\times$ 2
LLaMA-3.1-405B $\checkmark$ 3
GPT-4o $\times$ 4
GPT-4-Turbo $\times$ 5
LLaMA-2-70B $\checkmark$ 6
Table 1: Comparison of Different prominent Large Language Models
To further investigate the factors influencing explicit and implicit biases of LLMs, this study selects the LLaMA series of open-source LLMs as experimental subjects. These LLMs not only provide transparent information regarding their pre-training data scale and parameters but also offer both pre-trained base models and instruction-tuned versions, providing ideal conditions for studying the independent effects and interactions of different factors.
In examining the impact of training data scale and model parameters on both types of bias, this paper selected the pre-trained versions of LLaMA-2, LLaMA-3, LLaMA-3.1, and LLaMA-3.2 series models. These LLMs range in parameter size from 1B to 405B, with training data spanning from 1T tokens to over 15T tokens, providing extensive experimental comparison conditions. The detailed information about these selected LLMs is shown in Table 2
Model Parameters Pre-trained Tokens
LLaMA-2 7B 2T
13B 2T
70B 2T
LLaMA-3 8B 15T
70B 15T
LLaMA-3.1 405B 15T
LLaMA-3.2 1B 9T
3B 9T
Table 2: Configurations of LLaMA model series.
In studying the effects of alignment training on explicit and implicit biases, this research focused primarily on the 1B and 3B models from the LLaMA-3.2 series. These LLMs, with their smaller parameter, have relatively lower computational resource requirements, making them more suitable for multiple rounds of preference alignment training experiments.
Specifically, this study employs Direct Preference Optimization (DPO)Rafailov et al. (2024) combined with Low-Rank Adaptation (LoRA)Hu et al. (2022), conducting human preference alignment training at various levels on Anthropic’s open-source harmless-base RLHF datasetBai et al. (2022a). Compared to Reinforcement Learning from Human Feedback (RLHF)Ouyang et al. (2022), Direct Preference Optimization (DPO) offers a more streamlined training process by eliminating the need for complex reward models and reinforcement learning procedures, resulting in significantly simplified implementation and reduced computational overhead. To comprehensively evaluate the impact of preference alignment training on model behavior, this study implements a fine-grained training step design, conducting evaluations at 100-step intervals until observing the near-elimination of one type of bias. The alignment training is performed on a single NVIDIA RTX 4090 GPU. To balance computational resources and training effectiveness, a gradient accumulation strategy is adopted: setting the batch size to 32, processing 4 samples per iteration with 8 gradient accumulation steps. The temperature coefficient $\beta$ in DPO is set to 0.1.
Model Age Disability Gender Career Gender Occup. Race Science
Imp. Exp. Imp. Exp. Imp. Exp. Imp. Exp. Imp. Exp. Imp. Exp.
Gemini-2-Flash 0.71 0.01 0.91 0.14 0.94 0.12 1.00 0.38 0.45 0.01 0.70 0.01
Claude-3.5-Sonnet 0.86 0.00 0.85 0.17 0.89 0.21 0.95 0.55 0.32 0.01 0.66 0.14
LLaMA-3.1-405B 0.71 0.07 0.67 0.17 0.78 0.38 0.89 0.53 0.46 0.01 0.59 0.10
GPT-4o 0.72 0.00 0.86 0.05 0.81 0.14 0.97 0.26 0.37 0.01 0.79 0.01
GPT-4-Turbo 0.58 0.00 0.61 0.03 0.74 0.04 0.95 0.20 0.27 0.00 0.43 0.00
LLaMA-2-70B 0.45 0.00 0.49 0.00 0.48 0.00 0.73 0.00 0.29 0.00 0.55 0.01
Table 3: Stereotypical Score (SC) in measures of explicit and implicit biases for every subject across LLMs. LLMs demonstrate notable inconsistency between implicit and explicit biases, with implicit biases exhibiting strong stereotyping while explicit biases showing only mild stereotyping
5 Results and Analysis
5.1 Explicit vs. Implicit Bias
Figure 2: Average Stereotype Scores: Comparing Explicit and Implicit Bias Across LLMs.
Table 3 presents detailed stereotype scores for explicit and implicit biases across six dimensions: age, disability, gender career, gender occupation, race, and science. Figure 2 visualizes the overall explicit and implicit bias performance across LLMs, ordered by model capability based on Chatbot Arena platform. The data reveals a consistent pattern across various LLMs: a notable inconsistency between implicit and explicit biases across different attribute dimensions. While implicit biases consistently show strong stereotypical tendencies, explicit biases remain relatively mild. This inconsistency is evident across multiple attributes. Furthermore, analysis of the relation between model capability and bias performance reveals a significant finding: model general capability positively correlates with the degree of implicit bias. Specifically, models ranking higher on the Chatbot Arena platform tend to exhibit stronger implicit stereotypical biases.
5.2 Analysis of Bias Factors
5.2.1 Impact of Data Count and Parameters
Figure 3: Impact of pre-training data size and model parameters on explicit and implicit biases. Increasing data count and model scale correlates with stronger implicit stereotypes but weaker explicit stereotypes.
The impact of training data size and model parameters on explicit and implicit biases are illustrated in Figure 3. The experimental results reveal distinctly different trends in explicit and implicit biases as model parameters and training data increase: explicit bias shows a significant decrease with increased training data and parameters, particularly pronounced in large-scale LLM (above 70B). When the parameter scale reaches 405B, explicit bias nearly vanishes (stereotype score of 0.03). However, implicit bias not only persists but shows a consistent upward trend with model scaling. In models with 405B parameters, implicit bias reaches a substantial stereotype score of approximately 0.72.
These findings reveal the persistent nature of implicit bias. Traditional optimization approaches of increasing model parameters and training data size, while effective in mitigating explicit bias, actually intensify implicit bias while improving model performance. This suggests that addressing implicit bias, as a more profound form of bias, may require more targeted solutions rather than merely relying on scaling model size or training data.
5.2.2 Impact of Alignment Steps
Figure 4: Impact of alignment training steps on LLaMA-3.2-1B Biases. Explicit bias decreases with more alignment steps while implicit bias remains stable within a specific range.
Figure 5: Impact of alignment training steps on LLaMA-3.2-3B biases. Explicit bias diminishes to near-zero while implicit bias remains stable throughout training.
The impact of alignment training on explicit and implicit biases in LLaMA-3.2-1B and LLaMA-3.2-3B are illustrated in Figures 4 and 5.
The experimental results reveal significant differences in how alignment training affects explicit and implicit biases. For explicit bias, alignment training demonstrates a clear inhibitory effect: after 300 training steps, explicit bias shows a rapid decline, eventually approaching zero. This phenomenon is verified in both LLMs, indicating that human preference alignment training effectively reduces explicit bias in LLMs. However, for implicit bias, the effectiveness of alignment training is notably limited. Implicit bias levels remain relatively stable throughout the training process for both 1B and 3B models, maintaining around 0.2. This "resistance to intervention" suggests that traditional human preference alignment methods may struggle to address and improve implicit bias in models. Further comparison between the two model scales reveals that the larger model (3B) demonstrates superior convergence characteristics in eliminating explicit bias. Its explicit bias curve shows significant improvement around 200 steps, while the 1B model requires a longer training period to achieve similar effects.
These findings provide crucial insights into the formation mechanisms of model bias: while alignment training effectively suppresses explicit bias, its intervention on implicit bias remains notably limited. This significant disparity not only reveals the deep complexity of model bias but also indicates the need for future research to develop more targeted approaches in addressing implicit bias.
6 Conclusion
This paper presents a systematic investigation of explicit and implicit biases in large language models(LLMs), conducting an in-depth comparative analysis of these two forms of bias. Initially, drawing from social psychology theories and methodologies, we propose a self-reflection-based approach to measure both explicit and implicit biases in LLMs. This method establishes explicit bias measurement through models’ self-reflection on implicit bias, enabling comparative analysis of both bias types within the same social groups. Experimental results across multiple LLMs and topics reveal significant inconsistencies between explicit and implicit biases: while LLMs demonstrate low stereotypical tendencies in explicit bias, they exhibit strong stereotypical patterns in implicit bias. Further investigation explores key factors influencing these biases, including training data size, model parameter count, and alignment training. The findings indicate that while increased training data and model parameters significantly reduce explicit bias, implicit bias shows an intensifying trend. Additionally, alignment training (such as RLHF and DPO), while effective in suppressing explicit bias, has limited impact on implicit bias, which maintains relatively stable intensity throughout the alignment process. These systematic experiments reveal the mechanisms behind the divergence between explicit and implicit biases in models, while demonstrating the notable persistence of implicit bias and the limitations of traditional bias mitigation methods in addressing implicit bias issues.